# GORAG: Graph-based Retrieval Augmented Generation for Dynamic Few-shot Text Classification

### v3

@@ 3.3. Part 2: Graph Retrieval
To begin with, GORAG extract keywords $\mathcal{V}_{test}^{t}$ for each query text $t\in\mathcal{T}_{test}$ in the same manner with Equation (2), then, $\mathcal{V}_{test}^{t}$ would be splitted into two subsets:
@@
To achieve the adaptive retrieval, GORAG trys to find the minimum cost spanning tree (MST) that contain all keyword nodes within $\mathcal{V}_{exist}^{t}$, and then retrieve all label nodes within the generated MST as candidate labels. The intuition behind this approach is that a MST spans the entire graph to cover all given nodes with the smallest possible spanning cost. Consequently, label nodes within the generated MST can be considered important for demonstrating the features of the given keyword node set. As the generation of an MST is a classical combinatorial optimization problem with an optimal solution determined solely by the set of given nodes. By generating the MST and retrieving all label nodes within it, we eliminate the need for any human-defined thresholds, achieving adaptive retrieval.
Table 1. Statistics of the WOS, Reuters, and IFS-Rel datasets, where we divide their original labels into multiple rounds. We achieve a balanced testing set on the Reuters and IFS-Rel datasets by assigning each label 10 and 40 testing samples, respectively.
@@
(10) $s.t.$ $v\in\mathcal{V}^{r}_{t},\forall v\in\mathcal{V}_{exist}^{t}$
Then, since the subgraph nodes $\mathcal{V}^{r}_{t}$ contains both keyword nodes and label nodes, we take the label nodes ${\mathcal{Y}}^{r}_{t}\in\mathcal{V}^{r}_{t}$ in the gernerated sub-graph as our target candidate nodes for the text $t$.
Since this problem is the NP-hard (Su et al., 2020), it is infeasible to obtain the optimal result in polynomial time. Therefore, to solve this problem, we propose a greedy algorithm based on the Melhorn’s algorithm (Mehlhorn, 1988a). Our algorithm generates the candidate label set ${\mathcal{Y}}^{r}_{t}\subseteq\mathcal{Y}^{r}$ for text $t$ at the $r$-th round. The pseudo code of our algorithm is shown in Appendix C, following the Melhorn’s algorithm, our algorithm can also achieve 2-approximate.
Firstly, we calculate the shortest path between each keyword node $v$ in $\mathcal{V}_{exist}^{t}$ to all other nodes in $\mathcal{G}^{r}$ by calculating the minimum spanning tree $MST$ of $\mathcal{G}^{r}$ Here, the keyword nodes $v$ in $\mathcal{V}_{exist}^{t}$ are served as terminal nodes that determines the final generated candidate labels w.r.t. to the weighted graph $\mathcal{G}^{r}$ Secondly, we create a new auxiliary graph $H$ where the edges represent the shortest paths between the closest terminal nodes. Thirdly, we construct the minimum spanning tree $MST^{\prime}$ of the auxiliary graph $H$ , and then add the shortest paths between each two nodes in $MST^{\prime}$ to the Steiner Tree $ST$. Lastly our candidate labels $\hat{\mathcal{Y}}^{r}_{t}$ for text $t$ are calculated as the interselect of $ST$ and all target labels $\mathcal{Y}^{r}$ at round $r$: $ST\cap\mathcal{Y}^{r}$. Please refer to Appendix C for more details.
3.4. Part 3: Text Classification and Online Index
In this subsection, we introduce GORAG’s classification part, where it futher enrich the weighted graph by adding query text information with the online indexing mechanism, addressing the narrow source issue. To begin with, we introduce how GORAG performs text classification based the LLM. Specifically, given each unlabled query text $t$, GORAG utilizes LLM to predict its class label $y^{*}$ by constructing an input prompt $c_{t}$ as follows:
@@ 3.4. Part 3: Text Classification and Online Index
$c_{t}$ can be considered as the text concatenation of the extracted keywords $\mathcal{K}^{t}$ from the text $t$, and the candidate labels $\hat{\mathcal{Y}}^{r}_{t}$ obtained by Algorithm 2. Also, the $\mathcal{K}_{\hat{\mathcal{Y}}^{r}_{t}}=\{\mathcal{K}_{y_{i}}\}_{i=1}^{|\hat{\mathcal{Y}}^{r}_{t}|}$ and $\mathcal{K}_{y_{i}}$ are the representative keywords of each label $y_{i}\in\hat{\mathcal{Y}}^{r}_{t}$.
@@
which were generated by LLM with the label’s semantic label name if avaiable. Next, with a classification instruction prompt $p_{classify}$.
(13) $\vspace{-5px}y^{*}_{t}=LLM(p_{classify},c_{t}),$
the LLM would try to select the best-suited label $y^{*}_{t}\in\mathcal{Y}^{t}$ to annotate the text $t$.
@@
(14) $\vspace{-5px}\mathcal{V}^{r}=\mathcal{V}^{r}\cup\mathcal{V}_{notexist}^{t},\ \mathcal{E}^{r}=\mathcal{E}^{r}\cup\mathcal{E}_{oi}^{t},$
where $\mathcal{E}_{oi}^{t}=\{e_{v,y_{t}^{*}}|v\in\mathcal{V}_{notexist}^{t}\}$ denotes the set of all newly assigned edges $e_{v,y}$ between keyword node $v$ and its predicted label $y_{t}^{*}$, for these newly assigned edges, their weight is calculated with the edge weighting mechanism illustrated in Equation (5).
This online indexing mechanism has two purposes. Firstly, it incrementally enriches the weighted graph $\mathcal{G}^{r}$ by incorporating new, previously unseen keywords, thus expanding the graph’s vocabulary and improving its adaptability. Secondly, by linking these new keywords to their predicted labels with weighted edges, this mechanism captures their relevance and context more effectively, enabling better utilization of the graph for future predictions.
@@ 4. Experiments
In this section, we present the experimental evaluation of our framework GORAG on the DFSTC task. We compare the performance of GORAG against six effective baselines spanning three technical categories, using three datasets with distinct characteristics. Specifically, we first outline the experimental setup in Section 4.1, including details on datasets, and evaluation metrics. Due to the space limitation, we put the details of baselines, and hyperparameter configurations in Appendix Section E. Next, we report the experimental results, focusing on both effectiveness and efficiency evaluations, in Section 4.2. Finally, we conduct an ablation study and a case study, presented in Section 4.5 and Appendix Section E.5.
@@ 4.1.2. Datasets
We select the Web-Of-Science (WOS) dataset (Kowsari et al., 2017) and the Reuters few-shot text classification dataset by (Bao et al., 2020) for evaluating GORAG’s performance. We chose these datasets because they have a considerable amount of classes and text tokens, which is complex and can reflect the real-world scenario. As shown in Table 1, texts from these datasets can contain multiple sentences and have average tokens ranging from 105 to 200 per text. The IFS-Rel dataset was extended from its original version in (Xia et al., 2021), which only has 32 labels in total and an average of 25 tokens per text.
Table 2. Classification accuracy on WOS, Reuters, and IFS-Rel dataset.
Dataset Category Model Round
R1 R2 R3 R4
Propositionizer (Chen et al., 2023) 0.1241 0.1306 0.1297 0.1074 0.1421 0.1303 0.1771 0.1645 0.1603 0.1611 0.1637 0.1656
Compression RAG LongLLMLingua (Jiang et al., 2024a) 0.3806 0.3823 0.3901 0.2155 0.2202 0.2198 0.1770 0.1567 0.1608 0.1468 0.1382 0.1493
GraphRAG (Edge et al., 2024) 0.3852 0.3897 0.3906 0.2213 0.2197 0.2219 0.1816 0.1770 0.1786 0.1641 0.1634 0.1625
LightRAG (Guo et al., 2024) 0.3930 0.3806 0.3815 0.2202 0.2216 0.2145 0.1743 0.1767 0.1799 0.1625 0.1626 0.1632
GORAG 0.4866 0.4990 0.5134 0.4790 0.5109 0.5284 0.4736 0.4996 0.5234 0.4380 0.4717 0.4929
Reuters NN-based Entailment (Xia et al., 2021) 0.0000 0.0000 0.0000 0.0000 0.0000 0.0000 0.0000 0.0000 0.0000 0.0000 0.0000 0.0000
Long Context RAG NaiveRAG (Gao et al., 2023) 0.0000 0.0000 0.0375 0.0000 0.0000 0.0000 0.0000 0.0000 0.0000 0.0000 0.0000 0.0000
Propositionizer (Chen et al., 2023) 0.0000 0.0000 0.0000 0.0000 0.0000 0.0000 0.0000 0.0000 0.0000 0.0000 0.0000 0.0000
Compression-RAG LongLLMLingua (Jiang et al., 2024a) 0.0000 0.0000 0.0000 0.0000 0.0000 0.0000 0.0000 0.0000 0.0000 0.0258 0.0065 0.0085
GraphRAG (Edge et al., 2024) 0.1375 0.1625 0.1000 0.0688 0.0813 0.0500 0.0375 0.0417 0.0417 0.0291 0.0375 0.0375
LightRAG (Guo et al., 2024) 0.0500 0.1375 0.1125 0.0250 0.0813 0.0563 0.0333 0.0333 0.0333 0.0125 0.0333 0.0208
GORAG 0.0875 0.0875 0.1000 0.1750 0.1688 0.1438 0.1667 0.1958 0.2167 0.1667 0.2516 0.2193
Long Context RAG NaiveRAG (Gao et al., 2023) 0.4844 0.4719 0.4828 0.3711 0.3555 0.3656 0.2708 0.2750 0.2625 0.2516 0.2465 0.2527
Propositionizer (Chen et al., 2023) 0.1719 0.1828 0.1938 0.1602 0.1727 0.1750 0.1792 0.1552 0.1770 0.1574 0.1688 0.1664
Compression-RAG LongLLMLingua (Jiang et al., 2024a) 0.4703 0.5203 0.4297 0.3458 0.3055 0.3733 0.2452 0.2206 0.2713 0.2154 0.2029 0.1997
GraphRAG (Edge et al., 2024) 0.4797 0.4656 0.4734 0.3672 0.3641 0.3641 0.2740 0.2776 0.2708 0.2453 0.2465 0.2426
LightRAG (Guo et al., 2024) 0.4484 0.4234 0.4656 0.3531 0.3477 0.3523 0.2760 0.2760 0.2682 0.2340 0.2367 0.2391
GORAG 0.4875 0.5266 0.5313 0.4195 0.3781 0.4374 0.2953 0.2974 0.3328 0.2973 0.2571 0.3034
Table 3. The size of the constructed weighted graph in each round.
Dataset Model Round
R1 R2 R3 R4 After R4
Node # Edge # Node # Edge # Node # Edge # Node # Edge # Node # Edge #
WOS GORAG offline 1,149 1,405 3,681 4,587 4,947 6,204 5,438 6,879 5,438 6,879
GORAG 1,149 1,405 12,108 12,245 25,357 26,290 34,521 35,919 44,283 45,973
Reuters GORAG offline 117 119 192 200 263 284 331 366 331 366
GORAG 117 119 584 594 1,013 1,046 1,342 1,405 1,533 1,600
IFS-Rel GORAG offline 189 213 242 275 372 469 454 594 454 594
GORAG 189 213 5,299 5,503 9,417 9,934 13,293 14,074 16,705 17,486
We split the WOS dataset into 4, 6 and 8 rounds, the Reuters and IFS-Rel datasets into 4 rounds, in order to maintain a comparable number of new label and testing data in each round. Due to the space limitation, we mainly display our experinents on the 4 rounds spit version of both datasets in Table 2, and the experiment result of 6 and 8 rounds are shown in Figure 5(a) and 5(b). Different from the WOS and IFS-Rel datasets, the labels of Reuters are in numeric formats, and their text-formed label names are not available. According to Table 2, this feature makes many baselines fail to carry out classification task on the Reuters dataset. Further statistics of these two dataset being splitted into 4 rounds are shown in Table 1.
@@ 4.1.3. Evaluation Metrics
In this paper, aligning with previous DFSTC models (Xia et al., 2021; Meng et al., 2020), we use classification accuracy as the main evaluation metric. Given that LLMs can generate arbitrary outputs that may not precisely match the provided labels, we consider a classification correct only if the LLM’s output exactly matches the ground-truth label name or label number, if the model cannot give out a class within the provided target label set, we denote this classification as a hallucination. We further provide the macro recall score and the hallucination rate on the most complicated and imbalanced test dataset WOS in Table 8 in Appendix.
@@ 4.1.4. Baselines
As shown in Table 2, we compare GORAG’s performance with 7 baselines from 3 technical categories for few-shot experiments, they are NN-based Entailment (Xia et al., 2021), Long Context-based NaiveRAG (Gao et al., 2023), Propositionizer (Chen et al., 2023), and Compression-based LongLLMLingua (Jiang et al., 2024a), GraphRAG (Edge et al., 2024), and LightRAG (Guo et al., 2024). Further illustrations of these chosen baselines are listed in subsection E.1 in Appendix. For GraphRAG and NaiveRAG, we use the implementation from (nan, [n.d.]) which optimizes their original code and achieves better time efficiency while not affect the performance; For all other baselines, we use their official implementations.
Also, long text inputs affect LLM performance and increase costs (Chen et al., 2023; Jiang et al., 2024a; Liu et al., 2024a). WOS and Reuters datasets, with lengthy texts and many labels, caused GPU out-of-memory (e.g., llama3-8B, 1-shot, A100-40GB). Thus, we compare open-source LLMs with GORAG in the 0-shot setting (Table 4).

### v4

@@ 3.3. Part 2: Graph Retrieval
To begin with, GORAG extracts keywords $\mathcal{V}_{test}^{t}$ for each query text $t\in\mathcal{T}_{test}$ in the same manner with Equation (2), then, $\mathcal{V}_{test}^{t}$ would be split into two subsets:
@@
To achieve the adaptive retrieval, GORAG tries to find the minimum cost spanning tree (MST) that contain all keyword nodes within $\mathcal{V}_{exist}^{t}$, and then retrieves all label nodes within the generated MST as candidate labels. The intuition behind this approach is that an MST spans the entire graph to cover all given nodes with the smallest possible spanning cost. Consequently, label nodes within the generated MST can be considered important for demonstrating the features of the given keyword node set. As the generation of an MST is a classical combinatorial optimization problem with an optimal solution determined solely by the set of given nodes. By generating the MST and retrieving all label nodes within it, we eliminate the need for any human-defined thresholds, achieving adaptive retrieval.
Table 1. Statistics of the tested datasets, where we divide their original labels into multiple rounds. We achieve a balanced testing set on the IFS-Rel datasets by assigning each label 40 testing samples.
@@
(9) $s.t.\$$v\in\mathcal{V}^{r}_{t},\forall v\in\mathcal{V}_{exist}^{t}$
Then, since the subgraph node $\mathcal{V}^{r}_{t}$ set contains both keyword nodes and label nodes, we define the candidate label set $\hat{\mathcal{Y}}^{r}_{t}\in\mathcal{V}^{r}_{t}$ as the label nodes in the generated sub-graph for the text $t$ at round $r$.
Due to space limitations, we put the proof of the Adaptive Candidate Label Generation Problem to be NP-hard to Appendix B.
As the problem is NP-hard, it is infeasible to obtain the optimal result in polynomial time. Therefore, to solve this problem, we propose a greedy algorithm based on the Melhorn’s algorithm (Mehlhorn, 1988). Our algorithm generates the candidate label set $\hat{\mathcal{Y}}^{r}_{t}\subseteq\mathcal{Y}^{r}$ for text $t$ at the $r$-th round. The pseudo code of our algorithm is shown in Algorithm 2 in Appendix C , following the Melhorn’s algorithm, our algorithm can also achieve 2-Approximate.
Firstly, we calculate the shortest path between each keyword node $v$ in $\mathcal{V}_{exist}^{t}$ to all other nodes in $\mathcal{G}^{r}$ by calculating the minimum spanning tree $MST$ of $\mathcal{G}^{r}$ Here, the keyword nodes $v$ in $\mathcal{V}_{exist}^{t}$ serve as terminal nodes that determines the final generated candidate labels w.r.t. to the weighted graph $\mathcal{G}^{r}$. Secondly, we create a new auxiliary graph $H$ where the edges represent the shortest paths between the closest terminal nodes. Thirdly, we construct the minimum spanning tree $MST^{\prime}$ of the auxiliary graph $H$ , and then add the shortest paths between each two nodes in $MST^{\prime}$ to the Steiner Tree $ST$. Lastly our candidate labels $\hat{\mathcal{Y}}^{r}_{t}$ for text $t$ are calculated as the intersection of $ST$ and all target labels $\mathcal{Y}^{r}$ at round $r$: $ST\cap\mathcal{Y}^{r}$.
3.4. Part 3: Classification & Online Indexing
In this subsection, we first introduce GORAG’s classification and online graph indexing procedure, where it classifies the texts and further enriches the weighted graph by adding query text information to the created graph, addressing the narrow source issue. To begin with, we introduce how GORAG performs text classification based the LLM. Specifically, given each unlabeled query text $t$, GORAG utilizes LLM to predict its class label $y^{*}$ by constructing an input prompt $c_{t}$ as follows:
@@ 3.4. Part 3: Classification & Online Indexing
$c_{t}$ can be considered as the text concatenation of the extracted keywords $\mathcal{K}^{t}$ from the text $t$, and the candidate labels $\hat{\mathcal{Y}}^{r}_{t}$ obtained by Algorithm 2 in Appendix C, we omit the more complex input structures as they rarely improve performance, but lead to higher costs (Zhou et al., 2023; Liu et al., 2024b). Also, the $\mathcal{K}_{\hat{\mathcal{Y}}^{r}_{t}}=\{\mathcal{K}_{y_{i}}\}_{i=1}^{|\hat{\mathcal{Y}}^{r}_{t}|}$ and $\mathcal{K}_{y_{i}}$ are the representative keywords of each label $y_{i}\in\hat{\mathcal{Y}}^{r}_{t}$.
@@
which were generated by LLM with the label’s semantic label name if available. Next, with a classification instruction prompt $p_{classify}$.
(12) $y^{*}_{t}=LLM(p_{classify},c_{t}),$
the LLM would try to select the best-suited label $y^{*}_{t}\in\mathcal{Y}^{r}$ to annotate the text $t$.
@@
(13) $\vskip-3.01125pt\mathcal{V}^{r}=\mathcal{V}^{r}\cup\mathcal{V}_{notexist}^{t},\ \mathcal{E}^{r}=\mathcal{E}^{r}\cup\mathcal{E}_{oi}^{t},$
where $\mathcal{E}_{oi}^{t}=\{e_{v,y_{t}^{*}}|v\in\mathcal{V}_{notexist}^{t}\}$ denotes the set of all newly assigned edges $e_{v,y}$ between keyword node $v$ and its predicted label $y_{t}^{*}$, for these newly assigned edges, their weight is calculated with the edge weighting mechanism illustrated in Equation (4). The edge weights can reflect the semantic relevance of keywords to labels, and reduce the noise introduced by inaccurate LLM classification.
@@ 4. Experiments
Table 2. Classification accuracy on CAD, WOS, and IFS-Rel benchmarks.
Dataset Category Model Round 1 Round 2 Round 3 Round 4
CAD NN-based Entailment (Xia et al., 2021) 0.2125 0.2000 0.1208 0.1700 0.1060 0.1634 0.1177 0.0836 0.0785 0.0857 0.0700 0.0529
Long Context RAG NaiveRAG (Gao et al., 2023) 0.1875 0.1375 0.0958 0.1500 0.1100 0.0766 0.0802 0.0563 0.0392 0.0600 0.0405 0.0281
LongLLMLingua (Jiang et al., 2024) 0.1667 0.2417 0.2417 0.1709 0.2167 0.2200 0.1370 0.1587 0.1552 0.1092 0.1359 0.1285
Graph-based RAG GraphRAG (Edge et al., 2024) 0.3958 0.3625 0.3292 0.3167 0.2933 0.2833 0.1587 0.1621 0.1706 0.1200 0.1236 0.1322
LightRAG (Guo et al., 2025) 0.3875 0.3583 0.3373 0.3133 0.2867 0.2867 0.1638 0.1502 0.1571 0.1175 0.1102 0.1188
HippoRAG2 (gutiérrez2025ragmemory) 0.3667 0.2250 0.1792 0.2953 0.2000 0.1458 0.1820 0.1032 0.0746 0.1435 0.0668 0.0535
GORAG 0.6125 0.6125 0.6125 0.5867 0.5833 0.5882 0.3993 0.4010 0.4001 0.3133 0.3146 03139
LongLLMLingua (Jiang et al., 2024) 0.3806 0.3823 0.3901 0.2155 0.2202 0.2198 0.1770 0.1567 0.1608 0.1468 0.1382 0.1493
Graph-based RAG GraphRAG (Edge et al., 2024) 0.3852 0.3897 0.3906 0.2213 0.2197 0.2219 0.1816 0.1770 0.1786 0.1641 0.1634 0.1625
LightRAG (Guo et al., 2025) 0.3930 0.3806 0.3815 0.2202 0.2216 0.2145 0.1743 0.1767 0.1799 0.1625 0.1626 0.1632
HippoRAG2 (gutiérrez2025ragmemory) 0.4133 0.4245 0.4208 0.2552 0.2656 0.2666 0.2205 0.2210 0.2242 0.1906 0.1860 0.1915
GORAG 0.4866 0.4990 0.5134 0.4790 0.5109 0.5284 0.4736 0.4996 0.5234 0.4420 0.4717 0.4929
Long Context RAG NaiveRAG (Gao et al., 2023) 0.7343 0.7344 0.7406 0.5394 0.5406 0.5469 0.4719 0.4719 0.4688 0.5233 0.4938 0.5000
LongLLMLingua (Jiang et al., 2024) 0.8750 0.8906 0.9219 0.5734 0.5922 0.5806 0.4703 0.4741 0.5225 0.5819 0.4903 0.5038
Graph-based RAG GraphRAG (Edge et al., 2024) 0.6406 0.7813 0.8125 0.4406 0.4688 0.4719 0.3828 0.3734 0.3141 0.4297 0.4422 0.3922
LightRAG (Guo et al., 2025) 0.7578 0.8281 0.8594 0.4906 0.5155 0.5156 0.3802 0.4010 0.4427 0.4531 0.4804 0.4610
HippoRAG2 (gutiérrez2025ragmemory) 0.7313 0.7375 0.7406 0.4875 0.5016 0.4938 0.4297 0.4328 0.4438 0.4984 0.5008 0.4953
GORAG 0.9688 0.9688 0.9688 0.5938 0.6172 0.6250 0.4844 0.4948 0.5344 0.6165 0.4988 0.5469
In this section, we evaluate GORAG’s performance on the DFSTC task. Section 4.1 introduces the experimental setup, while Section 4.2 presents results on effectiveness and efficiency. Section 4.3 covers ablation studies, and Section 4.4 discusses generalization. Due to space limitation, we put experiment details, such as hardware settings and case study in Appendix E.
@@ 4.1.2. Datasets
We select the CAD (Vidgen et al., 2021), WOS (Kowsari et al., 2017), and IFS-Rel (Xia et al., 2021) benchmark to test GORAG’s performance. The texts in CAD are social media posts that potentially have different kinds of hate speech and abuse. The texts in the IFS-Rel benchmark are segments from web news, and their classifications help detect misinformation. The texts from the WOS benchmark are abstracts or introductions from academic papers, as the WOS dataset contains a vast number of candidate labels and has a large word count for each text; these characteristics make it a challenging benchmark for text classification models. The statistics of these benchmarks are shown in Table 1.
Due to the space limitations, we mainly display our experinents on the 4 rounds split version of all datasets in Table 2, and further experiment results of 6 and 8 rounds are shown in Figure 4.
Ethical Disclaimer. This research involves sensitive content. All datasets are publicly available, and our study adheres to institutional ethical guidelines to ensure that no personally identifiable information is disclosed. Examples shown in this paper are for illustrative purposes only and do not reflect the authors’ views.
@@ 4.1.3. Evaluation Metrics
In this paper, aligning with previous DFSTC models (Xia et al., 2021; Meng et al., 2020), we use classification accuracy as the evaluation metric. Given that LLMs can generate arbitrary outputs that may not precisely match the provided labels, we consider a classification correct only if the LLM’s output exactly matches the ground-truth label name or label number.
@@ 4.1.4. Baselines
We compare six baselines from three types as follows. The details of selected baselines are in Appendix E.1.
  • One NN-based model: Entailment (Xia et al., 2021), which converts the task to binary classification, and creates entailment pairs with the labeled data for training.
  • Two Long Context RAG models: NaiveRAG (Gao et al., 2023) splits few-shot labeled data into chunks and retrieves the most query similar chunks to augment LLM classification. LongLLMLingua (Jiang et al., 2024) further compresses the split chunks of NaiveRAG by filtering out chunk tokens regarding their LLM generation perplexity.
  • Three Graph-based RAG models: GraphRAG (Edge et al., 2024) extracts entities and relations from few-shot training data, and uses LLM to summarize graph neighborhoods as communities for retrieval. LightRAG (Guo et al., 2025) drops the community-related procedure and solely retrieves the 1-hop neighborhood of the unlabeled text mapped entities on the graph. HippoRAG2 (gutiérrez2025ragmemory) retrieves the Top-k candidate labels and their training texts based on their Personalized PageRank (PPR) score on the graph.
Furthermore, as solely put the labeled text within the LLM context can greatly increase costs and affect LLM performance (Chen et al., 2023; Jiang et al., 2024; Liu et al., 2024b), we compare the performance of solely using LLM for classification with GORAG under 0-shot setting in Table 6.
4.1.5. Hyperparameter Settings
Following the exact setting in Entailment’s original paper (Xia et al., 2021), we train the Entailment model with the RoBERTa-large PLM for 5 epochs with a batch size of 16, a learning rate of $1\times 10^{-6}$, and the Adam optimizer (Kingma, 2014). For LongLLMLingua, we test the compression rate within 0.75,0.8,0.85{0.75,0.8,0.85} and select 0.8, as it achieves the best overall classification accuracy. For all RAG-based baselines, we set all their hyperparameter the same as in their original paper, if not further illustrated. For GraphRAG and LightRAG, we use their local search mode, as it achieves the highest classification accuracy on the most complex WOS dataset. For GraphRAG and NaiveRAG, we use the implementation from (20), which optimizes their original code and achieves better time efficiency while not affect the performance. For all other baselines, we use their official implementations.